# Moments and the Basic Equilibrium Equations

### v2

@@ 6. Generalized Dziobek Equations
Let $S$ be the $c\times c$ matrix such that $\mathcal{A}=\mathcal{S}\mathcal{W}^{T}$; notice that $\mathfrak{F}$ is a symmetric family if and only if $S$ is symmetric. So, if $\mathfrak{F}$ is symmetric, we can rewrite equation (6.1) as
@@
Equation (6.2) contains the classical Dziobek equations for central configurations of codimension one [4, 16, 29], so we call (6.2) the generalized Dziobek equations. Observe that, in order to apply theorem 6.1 to central configurations, we must take $\mathcal{F}=\mathcal{C}\mathcal{M}$, where $\mathcal{C}$ is the matrix whose $j^{\text{th}}$ column is the weight vector $C_{j}$ defined by (5.4), and $\mathcal{M}$ is the diagonal matrix with entries $m_{1},\ldots,m_{n}$ on the main diagonal. Remarks 3.3 and 3.4 discuss other possible applications of theorem 6.1.
@@
The number of parameters defining $S$ is $c+1\choose 2$}, which, according to formula (A.2.4) in appendix A, is precisely the number of constraints that must be imposed on the mutual distances of a configuration of $n$ points to guarantee it has a given codimension $c>0$. This coincidence is not fortuitous since, by theorem A.2.2, the vanishing of $c+1\choose 2$} Cayley-Menger determinants provide the necessary constraints. Dziobek’s approach via critical point theory then introduces $c+1\choose 2$} Lagrange multipliers. Dziobek also verified that the derivative of a Cayley-Menger determinant with respect to a squared mutual distance can be expressed simply as a constant times the product of a pair of oriented volumes of suitable subconfigurations [16, 21]. Aside from determining the exact value of the constant, this fact can be readily inferred from theorem 6.1. Our proof of theorem 6.1 was inspired by the elegant proof of proposition 8 in [4].
@@ A.1. Direct constraints
Notice the expression on the right-hand side of the above equation is the value of the left-hand side at the centroid of the rhombus. This observation led us to search for simpler alternatives to the current methods for obtaining constraints for mutual distances of configurations, specially to the Cayley-Menger determinant conditions to be soon discussed. We encountered the following corollary to the HLK theorem.
@@
Let $(X,w)$ and $(X^{\prime},w^{\prime})$ be weighted systems whose total weights are nonzero and equal, and whose barycenters coincide. The difference between the respective second moments has the same value at all points, namely, the value it has at the (common) barycenter.
@@ A.2. Determinantal constraints
Let us view $\mu_{0}$ as a linear functional on the weight space $\mathbb{W}$. Notice corollary 2.6 ensures that $\text{Ker}\,\mathcal{B}(X)\cap\mu_{0}^{-1}(0)\leq\mathbb{W}_{0}(X)$ for arbitrary $X$, with equality occurring if and only if $\mu_{2}=0$ for any basis of $\mathbb{W}_{0}(X)$, a condition which is fulfilled, as we just saw, if $X$ is co-spherical. It is noteworthy that the matrix $\mathcal{B}(X)$ plays a key role in the reduction of the $n$-body problem by rotations and translations [3, 28]. It turns out that, up to a multiplicative constant, matrix $\mathcal{B}(X)$ represents a bilinear form known as the relative configuration matrix of $X$, for it actually characterizes $X$ up to isometries [3, 28]. We shall consider the kernel of $\mathcal{B}(X)$, viewed as a bilinear form, in appendix B.
@@
Let $X$ be a configuration of dimension $d$. The mutual distances of $X$ are such that $c+1\choose 2$} functionally independent Cayley-Menger determinants of subconfigurations of $X$ vanish, where $c=\emph{codim}(X)$.
@@ Appendix B Extended Leibniz Identities
In this appendix we show that the kernel of $\mathcal{B}(X)$, viewed as a symmetric bilinear form on the space $\mathbb{D}=\mu_{0}^{-1}(0)$ is precisely $\mathbb{W}_{0}(X)$, the space of weight vectors with $\mu_{1}=\overrightarrow{O}$. This basic remark, which appears in [3], produces relations of interest to applications. Here we provide an independent proof, followed by an application to central configurations.
@@
Acknowledgements. I thank Alain Albouy for early discussions on the project that lead to this work, the participants of the conference Celestial Mechanics in Itajubá, organized by Luis Fernando Mello and Antonio Carlos Fernandes, and the participants of the Mathematical Physics Seminar at UFPE, Recife, for their comments and suggestions.

### v3

@@ 6. Generalized Dziobek Equations
Let $\mathcal{S}$ be the $c\times c$ matrix such that $\mathcal{A}=\mathcal{S}\mathcal{W}^{T}$; notice that $\mathfrak{F}$ is a symmetric family if and only if $\mathcal{S}$ is symmetric. So, if $\mathfrak{F}$ is symmetric, we can rewrite equation (6.1) as
@@
Equation (6.2) contains the classical Dziobek equations for central configurations of codimension one [4, 16, 29], so we call (6.2) the generalized Dziobek equations. Observe that, in order to apply theorem 6.1 to central configurations, we must take $\mathcal{F}=\mathcal{C}\mathcal{M}$, where $\mathcal{C}$ is the matrix whose $j^{\text{th}}$ column is the weight vector $C_{j}$ defined by (5.4), and $\mathcal{M}$ is the diagonal matrix with entries $m_{1},\ldots,m_{n}$ on the main diagonal. Remark 3.3 indicates other possible applications of theorem 6.1.
@@
The number of parameters defining $S$ is $c+1\choose 2$}, which, according to formula (A.2.4) in appendix A, is precisely the number of constraints that must be imposed on the mutual distances of a configuration of $n$ points to guarantee it has a given codimension $c>0$. This coincidence is not fortuitous since, by theorem A.2.2, the vanishing of $c+1\choose 2$} Cayley-Menger determinants provide the necessary constraints. Dziobek’s approach via critical point theory then introduces $c+1\choose 2$} Lagrange multipliers. Dziobek also verified that the derivative of a Cayley-Menger determinant with respect to a squared mutual distance can be expressed simply as a constant times the product of a pair of oriented volumes of suitable subconfigurations [16, 21]. Aside from determining the exact value of the multiplicative constant, this fact can be readily inferred from theorem 6.1. Our proof of theorem 6.1 was inspired by the elegant proof of proposition 8 in [4].
@@ A.1. Direct constraints
Notice the expression on the right-hand side of the above equation is the value of the left-hand side at the centroid of the rhombus. This observation led us to search for simpler alternatives to the current methods for obtaining constraints for mutual distances of configurations, specially to the Cayley-Menger determinant conditions soon to be discussed. We encountered the following corollary to the HLK theorem.
@@
Let $(X,w)$ and $(X^{\prime},w^{\prime})$ be weighted systems in an Euclidean affine space whose total weights are nonzero and equal, and whose barycenters coincide. The difference between the respective second moments has the same value at all points, namely, the value it has at the (common) barycenter.
@@ A.2. Determinantal constraints
Let us view $\mu_{0}$ as a linear functional on the weight space $\mathbb{W}$ introduced in section 4. Notice corollary 2.6 ensures that $\text{Ker}\,\mathcal{B}(X)\cap\mu_{0}^{-1}(0)\leq\mathbb{W}_{0}(X)$ for arbitrary $X$, with equality occurring if and only if $\mu_{2}=0$ for any basis of $\mathbb{W}_{0}(X)$, a condition which is fulfilled, as we just saw, if $X$ is co-spherical. It is noteworthy that the matrix $\mathcal{B}(X)$ plays a key role in the reduction of the $n$-body problem by rotations and translations [3, 28]. Indeed, up to a multiplicative constant, $\mathcal{B}(X)$ represents a bilinear form known as the relative configuration matrix of $X$, for it actually characterizes $X$ up to isometries [3, 28]. We shall consider the kernel of $\mathcal{B}(X)$, viewed as a bilinear form, in appendix B.
@@
Let $X$ be a configuration of dimension d<n1d<n-1. The mutual distances of $X$ are such that $c+1\choose 2$} functionally independent Cayley-Menger determinants of subconfigurations of $X$ vanish, where $c=\emph{codim}(X)$.
@@ Appendix B Extended Leibniz Identities
In this appendix we show that the kernel of the relative configuration matrix $\mathcal{B}(X)$, viewed as a symmetric bilinear form on the space $\mathbb{D}=\mu_{0}^{-1}(0)$ is precisely $\mathbb{W}_{0}(X)$, the space of weight vectors with $\mu_{1}=\overrightarrow{O}$. This basic remark, which appears in [3], produces relations of interest to applications. Here we provide an independent proof, followed by an application to central configurations.
@@
Acknowledgements. I wish to thank Alain Albouy for early discussions on the project that lead to this work, and for the comments and suggestions received from the participants of events organized by Luis Fernando Mello, Antonio Carlos Fernandes and Lei Zhao, and the Mathematical Physics Seminar in Recife. I specially thank Samuel Adrian Antz for sending a list of corrections to a previous version of the manuscript.